OpenStudy (anonymous):

A manufacturer of a soft drink advertises their orange soda as "naturally flavoured", although it contains only 5% orange juice. A new federal regulation stipulates that to be called "natural", a drink must contain at least 10% fruit juice. How much pure orange juice must this manufacturer add to 900gal of orange soda to conform to the new regulation?

OpenStudy (anonymous):

900*0.05(juice)= 45gal. there is already 45 gal in and you need to double the concentration so add 45 gal

OpenStudy (anonymous):

900\[\times\].05=45 gals 900\[\times\].1=90 gals 90-45=45gals more

OpenStudy (anonymous):

So you have 5% right now, you are saying just double that ?

OpenStudy (anonymous):

yeah